# Rational Tuning of LLM Cascades via Probabilistic Modeling

### v3

@@ Abstract
Understanding the reliability of large language models (LLMs) has recently garnered significant attention. Given LLMs’ propensity to hallucinate, as well as their high sensitivity to prompt design, it is already challenging to predict the performance of an individual LLM. However, the problem becomes more complex for compound LLM systems such as cascades, where in addition to each model’s standalone performance, we must understand how the error rates of different models interact. In this paper, we present a probabilistic model for the joint performance distribution of a sequence of LLMs, which enables a framework for rationally tuning the confidence thresholds of a LLM cascade using continuous optimization. Compared to selecting confidence thresholds using grid search, our parametric Markov-copula model significantly improves runtime scaling with respect to the length of the cascade and the desired resolution of the cost-error curve, turning them from intractable into low-order polynomial. In addition, the optimal thresholds computed using our continuous optimization-based algorithm increasingly outperform those found via grid search as cascade length grows, improving the area under the cost-error curve by 1.9% on average for cascades with $k\geq 3$ models. Overall, our Markov-copula model provides a rational basis for tuning LLM cascade performance and points to the potential of probabilistic methods in analyzing LLM systems.
@@ 1 Introduction
In this paper, we present a parametric probabilistic model for the joint distribution of the calibrated confidences of a sequence of $k$ LLMs, providing a rational basis for understanding the performance of LLM cascades. We focus on cascades whose constituent models are ordered by size, from smallest to largest. Our probabilistic model is based on a Markov factorization, leveraging the insight that LLMs similar in size are more predictive of each other’s confidence. After using logistic regression to calibrate each LLM’s confidence, we account for the pairwise interactions between subsequent LLMs’ error rates using bivariate copulas, providing a data-efficient model of cascade performance that performs well with $n\approx 300$ training examples across six benchmarks.
Our Markov-copula model makes it possible to tune the confidence thresholds of an LLM cascade using continuous optimization. Compared to selecting these thresholds via grid search, our algorithm significantly improves runtime scaling. For example, we reduce scaling with respect to the cascade length $k$ from exponential to low-order polynomial, making it much faster to tune longer cascades consisting of $k\geq 5$ models. In addition, the optimal thresholds computed using our continuous optimization-based algorithm increasingly outperform those found via grid search as cascade length grows. For cascades consisting of $k\geq 3$ models, we report an average $1.9\%$ decrease in the area under the cost-error curve across six benchmarks.
@@
  • Building on our Markov-copula model, we develop an algorithm for tuning the confidence thresholds of an LLM cascade using continuous optimization. We demonstrate that our method significantly improves the computational complexity of finding optimal confidence thresholds, turning the dependencies on cascade length and the desired resolution of the cost-error curve from intractable and high-order polynomial into low-order polynomial and linear, respectively. In addition, we show that our method finds higher-quality optimal thresholds as cascade length grows, reporting a 1.9% average decrease in the area under the cost-error curve for cascades consisting of $k\geq 3$ models.
@@ 3.3 Tuning the Confidence Thresholds
The purpose of the Markov model (6) is to obtain optimal cost-error tradeoffs for an LLM cascade $C$ by tuning the confidence thresholds. We formulate the optimization problem
$\boldsymbol{\theta}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}~{}(1-\mathbb{P}_{\boldsymbol{\theta}}(\text{Correct}))+\lambda~{}\mathbb{E}_{\boldsymbol{\theta}}[\text{Cost}],$ (11)
where $\boldsymbol{\theta}\in\mathbb{R}^{k-1}$ denotes the confidence thresholds $(\phi_{1},...,\phi_{k-1})$. The Lagrange multiplier $\lambda\geq 0$ indicates the user’s cost sensitivity. Setting $\lambda=0$ means that cost is irrelevant, whereas $\lambda>0$ penalizes the use of expensive models. To compute the efficient frontier of optimal $(\mathbb{P}(\text{Correct}),\mathbb{E}[\text{Cost}])$ tuples, we solve (11) for different values of the cost sensitivity $\lambda$.
@@
for some $r>0$, until the cost constraint is stringent enough to make the expected cost equal to the least expensive model’s expected cost. Typically, setting $r$ between 0.25 and 1 performs well. For any potential gaps in coverage, we adaptively interpolate the optimal thresholds. Specifically, if $\lambda^{(i)}<\lambda^{(i+1)}$ yield optimal thresholds $\boldsymbol{\theta}^{(i)}$ and $\boldsymbol{\theta}^{(i+1)}$ and the gap $|\boldsymbol{\theta}_{j}^{(i+1)}-\boldsymbol{\theta}_{j}^{(i)}|=|\phi_{j}^{(i+1)}-\phi_{j}^{(i)}|$ for any individual threshold exceeds probability mass $q$ based on the distribution of the calibrated confidence $\Phi_{j}$, we insert
$\boldsymbol{\theta}^{(i+1/2)}=(\boldsymbol{\theta}^{(i)}+\boldsymbol{\theta}^{(i+1)})/2$ (13)
into the list of optimal thresholds between $\boldsymbol{\theta}^{(i)}$ and $\boldsymbol{\theta}^{(i+1)}$. We repeat the infilling procedure (13) until no gaps remain at level $q$. We have found $q<0.2$ to perform well.
Efficient computation and optimization of the objective: solving the minimization problem (11) requires computing a cascade’s probability of correctness and expected cost for candidate confidence thresholds $\boldsymbol{\theta}=(\phi_{1},...,\phi_{k-1})\in\mathbb{R}^{k-1}$. To compute these quantities, we rely on the decompositions (14) and (15) presented in
@@ 4.1 Methodology
Benchmarks: we evaluate our probabilistic model and the cost-error curves of LLM cascades on six language modeling benchmarks including MMLU (Hendrycks et al.,, 2021); MedMCQA (Pal et al.,, 2022); TriviaQA (Joshi et al.,, 2017); XSum (Narayan et al.,, 2018); GSM8K (Cobbe et al.,, 2021); and TruthfulQA (Lin et al., 2022b, ). These tasks include general-purpose knowledge and reasoning, domain-specific QA, open-domain QA, summarization, mathematical reasoning, and truthfulness in the face of adversarially chosen questions.
@@
Importantly, we run each benchmark in a zero-shot manner, since we believe this setting faithfully reflects off-the-shelf use of LLMs in practice. Appendix C gives the prompt templates we used for each benchmark. To conveniently transform and calibrate the raw confidence scores, track the numbers of input and output tokens, and monitor cost, we ran our evaluations using a preliminary version of the niagara Python package for LLM cascading. Code for reproducing the results of the paper is available on GitHub. [^3] Code for reproducing the results of the paper is available at github.com/mzelling/rational-llm-cascades.
@@ 4.2 Performance Summary
MMLU MedMCQA TriviaQA XSum GSM8K TruthfulQA Model %ECE %ECE-${}_{\text{-TF}}$$\boldsymbol{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\boldsymbol{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\boldsymbol{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\boldsymbol{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\boldsymbol{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\boldsymbol{\%\Delta}$ llama3.2-1b 3.8 6.1 -37.7 8.7 9.8 -11.2 5.8 5.8 0.0 2.9 2.7 7.4 13.1 13.2 -0.8 4.3 4.3 0.0 llama3.2-3b 4.0 7.4 -45.9 6.8 10.0 -32.0 4.5 14.9 -69.8 3.6 3.7 -2.7 9.5 9.5 0.0 7.5 6.4 17.2 llama3.1-8b 4.1 7.0 -41.4 9.4 14.1 -33.1 6.2 15.1 -58.9 3.5 9.7 -63.9 4.7 5.1 -7.8 7.3 7.3 0.0 llama3.1-70b 2.4 7.9 -69.6 9.9 12.5 -20.8 2.3 5.1 -54.9 6.0 10.4 -42.3 2.9 4.5 -35.6 5.7 5.3 7.5 llama3.1-405b 2.9 10.4 -72.1 10.8 14.2 -24.0 3.0 4.9 -38.8 5.4 10.2 -47.1 1.9 3.6 -47.2 5.6 10.8 -48.1 qwen2.5-32b-c 5.3 13.9 -61.9 6.2 14.5 -57.2 10.0 15.7 -36.3 4.3 10.2 -57.8 3.2 4.7 -31.9 5.9 10.6 -44.3 qwen2.5-72b 4.9 10.9 -55.0 7.0 16.1 -56.5 4.0 9.4 -57.4 2.2 3.2 -31.3 1.2 4.9 -75.5 7.5 9.3 -19.4 gpt-4o-mini 4.7 14.8 -68.2 5.3 15.5 -65.8 1.4 3.8 -63.2 2.6 2.4 8.3 3.5 6.1 -42.6 5.3 5.8 -8.6 gpt-4o 4.8 11.2 -57.1 3.1 12.6 -75.4 2.1 3.6 -41.7 0.7 1.0 -29.6 2.1 4.6 -54.3 3.8 11.4 -66.7 Average 5.0 7.9 -36.7 7.6 9.8 -22.5 4.5 8.2 -45.0 3.1 4.3 -28.2 6.2 7.1 -12.7 5.1 6.7 -23.9
@@ 4.3.1 Verifying the Markov Assumption
These findings suggest that a cascade composed only of Llama models (1B-405B) satisfies the Markov assumption more exactly. Consider Figure 1a as an example. For the Llama cascade, Kendall’s $\tau$ is highest near the heatmap’s diagonal, suggesting a Markov property. By contrast, the mixed cascade composed of Llama, GPT, and Qwen models shows a more haphazard pattern. For example, the rank correlation between GPT-4o Mini and GPT-4o ($\tau=0.55$) is higher than that between GPT-4o and Llama3 405B ($\tau=0.54$), even though the latter pair of models are more similar in size. Similarly, Llama3 405B is more strongly correlated with Llama3 70B ($\tau=0.58$) than with Qwen2.5 72B ($\tau=0.46$), even though the latter models are of near-identical size. These examples highlight that, in order for the Markov property to hold based on model size, it seems important that models share the same architectural family.
@@ 4.3.3 Testing the Discrete-Continuous Marginal Confidence Distributions
$\sqrt{\text{CvM}}=\sqrt{\int(\hat{F}_{n}^{\text{test}}(x)-F(x|\boldsymbol{\hat{\theta}}))^{2}\text{~{}d}F(x|\boldsymbol{\hat{\theta}})},$ (20)
where $\hat{F}_{n}^{\text{test}}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\Phi(x_{i})}$ is the empirical distribution of the calibrated confidence on the test data, and $F(\cdot|\boldsymbol{\theta})$ is our marginal distribution model (7) with $\boldsymbol{\theta}=(\phi_{\text{min}},\phi_{\text{max}},w_{\text{min}},w_{\text{max}},\pi,\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})$. In Tables 5 and 6, we report (20) both for $\boldsymbol{\hat{\theta}}$ estimated from the training data ($\sqrt{\text{CvM}}$), and for $\boldsymbol{\hat{\theta}}$ re-fitted on the test data ($\sqrt{\text{CvM}_{\text{r}}}$). The reason we report $\sqrt{\text{CvM}_{\text{r}}}$ is to evaluate whether deficiencies in the fit arise from a bias problem, rather than a variance problem. To compute $p$ values for (20), we use parametric bootstrapping with $B=1000$ samples.
@@ 4.4 Rational Tuning of Confidence Thresholds
Figure 5: Shows the reduction in the area under the cost-error curve (AUC) on the test set when using our continuous-optimization-based method for selecting confidence thresholds, compared to grid search, as a function of cascade length. For longer cascades, our method outperforms by larger margins. Error bars show ±$\pm 1\sigma$ of the mean percentage change, and filled markers indicate statistical significance.
Benchmark AUC $\boldsymbol{\downarrow}$$\boldsymbol{\Delta}$ $\boldsymbol{p}$
Continuous Grid
MMLU 0.288 0.288 -0.001 $3.0\times 10^{-1}$
MedMCQA 0.381 0.384 -0.776 $\boldsymbol{4.7\times 10^{-3}}$
TriviaQA 0.182 0.183 -0.705 $\boldsymbol{7.6\times 10^{-3}}$
XSum 0.408 0.414 -1.485 $\boldsymbol{8.2\times 10^{-7}}$
GSM8K 0.159 0.163 -2.482 $\boldsymbol{3.7\times 10^{-4}}$
TruthfulQA 0.437 0.438 -0.288 $\boldsymbol{1.4\times 10^{-2}}$
Average 0.309 0.312 -0.956 -

### v4

@@ Abstract
Understanding the reliability of large language models (LLMs) has recently garnered significant attention. Given LLMs’ propensity to hallucinate, as well as their high sensitivity to prompt design, it is already challenging to predict the performance of an individual LLM. However, the problem becomes more complex for compound LLM systems such as cascades, where in addition to each model’s standalone performance, we must understand how the error rates of different models interact. In this paper, we present a probabilistic model for the joint performance distribution of a sequence of LLMs, which enables a framework for rationally tuning the confidence thresholds of a LLM cascade using continuous optimization. Compared to selecting confidence thresholds using Bayesian optimization, our parametric Markov-copula model yields more favorable error-cost trade-offs, improving the area under the error-cost curve by 4.3% on average for cascades with $k\geq 3$ models. In the low-sample regime with $n\leq 30$ training examples, the performance improvement widens to 10.2%, suggesting that our framework’s inductive assumptions about the interactions between the error rates of different LLMs enhance sample efficiency. Overall, our Markov-copula model provides a rational basis for tuning LLM cascade performance and points to the potential of probabilistic methods in analyzing systems of LLMs.
@@ 1 Introduction
In this paper, we present a parametric probabilistic model for the joint distribution of the calibrated confidences of a sequence of $k$ LLMs, providing a rational basis for understanding the performance of LLM cascades. We focus on cascades whose constituent models are ordered by size, from smallest to largest. Our probabilistic model is based on a Markov factorization, leveraging the insight that LLMs similar in size are more predictive of each other’s confidence. After using logistic regression to calibrate each LLM’s confidence, we account for the pairwise interactions between subsequent LLMs’ error rates using bivariate copulas, providing a data-efficient model of cascade performance that performs well with as few as $n\leq 30$ training examples across six benchmarks.
Our Markov-copula model makes it possible to tune the confidence thresholds of an LLM cascade using continuous optimization.
Compared to selecting these thresholds via Bayesian optimization, our Rational Tuning framework yields increasingly better error-cost trade-offs as cascade length grows. For cascades with $k\geq 3$ models, our method improves the area under the error-cost curve by 4.3% on average. Compared to high-resolution grid search, the improvement is 2.0%. At the same time, our algorithm significantly improves runtime scaling compared to grid search. For example, we reduce scaling with respect to the cascade length $k$ from exponential to low-order polynomial, making it much faster to tune longer cascades consisting of $k\geq 5$ models.
@@
  • Building on our Markov-copula model, we develop an algorithm for tuning the confidence thresholds of an LLM cascade using continuous optimization, based on an analytic probabilistic model. We demonstrate that as cascade length grows, our method increasingly outperforms the error-cost trade-offs obtained with Bayesian optimization and high-resolution grid search baselines. In addition, relative to grid search our method significantly improves the computational complexity of finding optimal confidence thresholds, turning the dependencies on cascade length and the desired resolution of the error-cost curve from intractable and high-order polynomial into low-order polynomial and linear, respectively.
@@ 3.3 Tuning the Confidence Thresholds
The purpose of the Markov model (6) is to obtain optimal error-cost tradeoffs for an LLM cascade $C$ by tuning the confidence thresholds. We formulate the optimization problem
$\bm{\theta}^{*}=\operatorname*{arg\,min}_{\bm{\theta}}~{}(1-\mathbb{P}_{\bm{\theta}}(\text{Correct}))+\lambda~{}\mathbb{E}_{\bm{\theta}}[\text{Cost}],$ (11)
where $\bm{\theta}\in\mathbb{R}^{k-1}$ denotes the confidence thresholds $(\phi_{1},...,\phi_{k-1})$. The Lagrange multiplier $\lambda\geq 0$ indicates the user’s cost sensitivity. Setting $\lambda=0$ means that cost is irrelevant, whereas $\lambda>0$ penalizes the use of expensive models. To compute the efficient frontier of optimal $(\mathbb{P}(\text{Correct}),\mathbb{E}[\text{Cost}])$ tuples, we solve (11) for different values of the cost sensitivity $\lambda$.
@@
for some $r>0$, until the cost constraint is stringent enough to make the expected cost equal to the least expensive model’s expected cost. Typically, setting $r$ between 0.25 and 1 performs well. For any potential gaps in coverage, we adaptively interpolate the optimal thresholds. Specifically, if $\lambda^{(i)}<\lambda^{(i+1)}$ yield optimal thresholds $\bm{\theta}^{(i)}$ and $\bm{\theta}^{(i+1)}$ and the gap $|\bm{\theta}_{j}^{(i+1)}-\bm{\theta}_{j}^{(i)}|=|\phi_{j}^{(i+1)}-\phi_{j}^{(i)}|$ for any individual threshold exceeds probability mass $q$ based on the distribution of the calibrated confidence $\Phi_{j}$, we insert
$\bm{\theta}^{(i+1/2)}=(\bm{\theta}^{(i)}+\bm{\theta}^{(i+1)})/2$ (13)
into the list of optimal thresholds between $\bm{\theta}^{(i)}$ and $\bm{\theta}^{(i+1)}$. We repeat the infilling procedure (13) until no gaps remain at level $q$. We have found $q<0.2$ to perform well.
Efficient computation and optimization of the objective: solving the minimization problem (11) requires computing a cascade’s probability of correctness and expected cost for candidate confidence thresholds $\bm{\theta}=(\phi_{1},...,\phi_{k-1})\in\mathbb{R}^{k-1}$. To compute these quantities, we rely on the decompositions (14) and (15) presented in
@@ 4.1 Methodology
Benchmarks: we evaluate our probabilistic model and the error-cost curves of LLM cascades on six language modeling benchmarks including MMLU (Hendrycks et al.,, 2021); MedMCQA (Pal et al.,, 2022); TriviaQA (Joshi et al.,, 2017); XSum (Narayan et al.,, 2018); GSM8K (Cobbe et al.,, 2021); and TruthfulQA (Lin et al., 2022b, ). These tasks include general-purpose knowledge and reasoning, domain-specific QA, open-domain QA, summarization, mathematical reasoning, and truthfulness in the face of adversarially chosen questions.
@@
Importantly, we run each benchmark in a zero-shot manner, since we believe this setting faithfully reflects off-the-shelf use of LLMs in practice. Appendix C gives the prompt templates we used for each benchmark. To conveniently transform and calibrate the raw confidence scores, track the numbers of input and output tokens, and monitor cost, we ran our evaluations using a preliminary version of the niagara Python package for LLM cascading. Code for reproducing the results of the paper is available on GitHub. [^3] Code for reproducing the results of the paper is available at github.com/mzelling/rational-llm-cascades..
@@ 4.2 Performance Summary
MMLU MedMCQA TriviaQA XSum GSM8K TruthfulQA Model %ECE %ECE-${}_{\text{-TF}}$$\bm{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\bm{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\bm{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\bm{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\bm{\%\Delta}$ECE %ECE-${}_{\text{-TF}}$$\bm{\%\Delta}$ llama3.2-1b 3.8 6.1 -37.7 8.7 9.8 -11.2 5.8 5.8 0.0 2.9 2.7 7.4 13.1 13.2 -0.8 4.3 4.3 0.0 llama3.2-3b 4.0 7.4 -45.9 6.8 10.0 -32.0 4.5 14.9 -69.8 3.6 3.7 -2.7 9.5 9.5 0.0 7.5 6.4 17.2 llama3.1-8b 4.1 7.0 -41.4 9.4 14.1 -33.1 6.2 15.1 -58.9 3.5 9.7 -63.9 4.7 5.1 -7.8 7.3 7.3 0.0 llama3.1-70b 2.4 7.9 -69.6 9.9 12.5 -20.8 2.3 5.1 -54.9 6.0 10.4 -42.3 2.9 4.5 -35.6 5.7 5.3 7.5 llama3.1-405b 2.9 10.4 -72.1 10.8 14.2 -24.0 3.0 4.9 -38.8 5.4 10.2 -47.1 1.9 3.6 -47.2 5.6 10.8 -48.1 qwen2.5-32b-c 5.3 13.9 -61.9 6.2 14.5 -57.2 10.0 15.7 -36.3 4.3 10.2 -57.8 3.2 4.7 -31.9 5.9 10.6 -44.3 qwen2.5-72b 4.9 10.9 -55.0 7.0 16.1 -56.5 4.0 9.4 -57.4 2.2 3.2 -31.3 1.2 4.9 -75.5 7.5 9.3 -19.4 gpt-4o-mini 4.7 14.8 -68.2 5.3 15.5 -65.8 1.4 3.8 -63.2 2.6 2.4 8.3 3.5 6.1 -42.6 5.3 5.8 -8.6 gpt-4o 4.8 11.2 -57.1 3.1 12.6 -75.4 2.1 3.6 -41.7 0.7 1.0 -29.6 2.1 4.6 -54.3 3.8 11.4 -66.7 Average 5.0 7.9 -36.7 7.6 9.8 -22.5 4.5 8.2 -45.0 3.1 4.3 -28.2 6.2 7.1 -12.7 5.1 6.7 -23.9
@@ 4.3.1 Verifying the Markov Assumption
Our findings suggest that a cascade composed only of Llama models (1B-405B) satisfies the Markov assumption more exactly. Consider Figure 1a as an example. For the Llama cascade, Kendall’s $\tau$ is highest near the heatmap’s diagonal, suggesting a Markov property. By contrast, the mixed cascade composed of Llama, GPT, and Qwen models shows a more haphazard pattern. For example, the rank correlation between GPT-4o Mini and GPT-4o ($\tau=0.55$) is higher than that between GPT-4o and Llama3 405B ($\tau=0.54$), even though the latter pair of models are more similar in size. Similarly, Llama3 405B is more strongly correlated with Llama3 70B ($\tau=0.58$) than with Qwen2.5 72B ($\tau=0.46$), even though the latter models are of near-identical size. These examples highlight that, in order for the Markov property to hold based on model size, it seems important that models share the same architectural family.
In Appendix F, we further verify the rank correlation patterns between different LLMs by recomputing the rank correlations only on those queries where both models answer correctly or incorrectly.
@@ 4.3.3 Testing the Discrete-Continuous Marginal Confidence Distributions
$\sqrt{\text{CvM}}=\sqrt{\int(\hat{F}_{n}^{\text{test}}(x)-F(x|\bm{\hat{\theta}}))^{2}\text{~{}d}F(x|\bm{\hat{\theta}})},$ (20)
where $\hat{F}_{n}^{\text{test}}=\frac{1}{n}\sum_{i=1}^{n}\delta_{\Phi(x_{i})}$ is the empirical distribution of the calibrated confidence on the test data, and $F(\cdot|\bm{\theta})$ is our marginal distribution model (7) with $\bm{\theta}=(\phi_{\text{min}},\phi_{\text{max}},w_{\text{min}},w_{\text{max}},\pi,\alpha_{1},\beta_{1},\alpha_{2},\beta_{2})$. In Tables 5 and 6, we report (20) both for $\bm{\hat{\theta}}$ estimated from the training data ($\sqrt{\text{CvM}}$), and for $\bm{\hat{\theta}}$ re-fitted on the test data ($\sqrt{\text{CvM}_{\text{r}}}$). The reason we report $\sqrt{\text{CvM}_{\text{r}}}$ is to evaluate whether deficiencies in the fit arise from a bias problem, rather than a variance problem. To compute $p$ values for (20), we use parametric bootstrapping with $B=1000$ samples.
@@ 4.4 Rational Tuning of Confidence Thresholds
(a) Error-cost curves
(b) Area under the error-cost curve
Figure 5: Performance evaluation via the area under the error-cost curve (AUC). (a) Error-cost curves computed on the MedMCQA test set for the Llama3 3B $\rightarrow$ 8B $\rightarrow$ 70B $\rightarrow$ 405B cascade. (b) Illustration of the area under the error-cost curve (AUC).
(a) Relative to Bayesian optimization
(b) Relative to grid search
@@
Figure 6: Reduction in the area under the error-cost curve (AUC) on the test set when using our Rational Tuning framework to select confidence thresholds, as a function of cascade length. In (a), we compare against a Bayesian optimization baseline, while in (b) we compare against high-resolution grid search. For longer cascades, our method outperforms both baselines by larger margins. Error bars show ±$\pm 1\sigma$ of the mean percentage change, and filled markers indicate statistical significance.
Benchmark AUC $\bm{\downarrow}$ RT (Ours) vs GS RT (Ours) vs BO
RT (Ours) GS BO %$\%\Delta_{\mathrm{GS}}\downarrow$ $p_{\mathrm{GS}}$$\%\Delta_{\mathrm{BO}}\downarrow$ $p_{\mathrm{BO}}$
MMLU 0.288 0.288 0.294 0.02 $2.8\times 10^{-1}$2.30 $\bm{4.3\times 10^{-3}}$
MedMCQA $\bm{0.381}$ 0.384 0.389 -0.79 $\bm{5.5\times 10^{-3}}$2.14 $\bm{3.0\times 10^{-2}}$
TriviaQA $\bm{0.181}$ 0.182 0.183 -0.74 $\bm{6.0\times 10^{-3}}$1.58 $\bm{1.0\times 10^{-2}}$
XSum $\bm{0.409}$ 0.414 0.416 -1.48 $\bm{2.0\times 10^{-6}}$1.92 $\bm{8.9\times 10^{-4}}$
GSM8K $\bm{0.158}$ 0.162 0.160 -2.68 $\bm{1.0\times 10^{-5}}$1.15 $\bm{1.6\times 10^{-2}}$
TruthfulQA $\bm{0.436}$ 0.437 0.438 -0.26 $\bm{3.8\times 10^{-2}}$0.46 $7.9\times 10^{-2}$
Average $\bm{0.309}$ 0.311 0.313 -0.99 -1.60